OpenStudy (nuccioreggie):

Two lines, E and F, are represented by the equations given below. Line E: 5x + 5y = 40 Line F: x + y = 8 Which statement is true about the solution to the set of equations? It is (40, 8). It is (8, 40). There is no solution. There are infinitely many solutions.

OpenStudy (whovianchick):

yooooooo wassup

OpenStudy (whovianchick):

im here XD

OpenStudy (nuccioreggie):

wassup lol

OpenStudy (nuccioreggie):

thanks lol this is math are you good with math

OpenStudy (whovianchick):

Um yeah I was the person who helped you all Friday or something with your 9 questions

OpenStudy (nuccioreggie):

i know who you are i gave you all those medals lol

OpenStudy (whovianchick):

x+y=8 -x -x y= 8-x I set this as y equals so that I can plug it into the other equation 5x+5y=40 5x+5(8-x)=40 We've established that y= 8-x, so you can put 8-x in place of y 5x+40-5x=40 Distribute 0x+40=40 Combine like terms 40=40 There are infinite solutions. These lines are the exact same line. Watch this x+y=8 5(x) + 5(y) = 5(8) 5x+5y=40 See?

OpenStudy (nuccioreggie):

so its C

OpenStudy (whovianchick):

The reason they have infinite solutions is because they intersect, or touch in a million places.

OpenStudy (whovianchick):

no it's D

OpenStudy (whovianchick):

There are infinite solutions

OpenStudy (whovianchick):



OpenStudy (whovianchick):

Since they're the same line, they have infinitely many intersections, or solutions

OpenStudy (whovianchick):